# Joint Reconstruction of Spatially-Coherent and Realistic Clothed Humans and Objects from a Single Image

### v1

@@ 9.3 Inpainting module
SCHOR (ours) 14.23 17.15 .8902 .8241 48.84
Table 6: Quantitative evaluation of changing the diffusion module of SCHOR. The human-object shapes reconstructed with the attention-based neural implicit module of SCHOR are used for evaluation.
@@
In scenes containing both humans and objects, the object often occlude parts of the human body. When only partial regions of the body are visible, the network fails to reconstruct the occluded regions. This limitation is shown in Fig. 10, where SCHOR is applied without the inpainting module to generate a full-body human image. Instead, the partial human image is directly input into the attention-based neural implicit model, which fails to reconstruct these occluded regions, producing holes in the human shapes. Inpainting the missing body regions in the input image is crucial to reconstruct realistic human-object shapes.
SCHOR addresses the challenge of human occlusion by leveraging the generative capability of diffusion models to inpaint the occluded body regions. A fine-tuning strategy is adopted to optimize the cross-attention layers of a pre-trained diffusion model, conditioning it on the mask of the body regions requiring inpainting. In this section, we explore different solutions for inpainting, including MAT [15], a transformer-based model for large hole inpainting, conditioning the diffusion model with the SMPL [27] normal map (Normal), and conditioning with both the mask and the normal map (Both). The effect of using these approaches for the inpainting module of SCHOR is quantitatively evaluated by estimating the implicit representation of human-object shapes using their generated full-body human image (Tab. 6). Examples of human-object reconstructions using these methods for inpainting are shown in Fig. 12 while Fig. 11 illustrates the input images shown in the paper and generated using the discussed inpainting methods.
@@
Note that we do not quantitatively evaluate the generated images against ground truth, since the main goal of SCHOR is 3D reconstruction rather than inpainting.
Figure 12: Human-object shapes reconstructed by changing the diffusion module of SCHOR. Front view with and without the object and side views without the object are shown from synHOR in the top row, from BEHAVE [2] in the bottom.
@@ 10 Additional visual comparisons
In this section, we present additional visual results of human-object shapes reconstructed using SCHOR, as well as comparisons with related methods that aim to reconstruct high-quality 3D humans (PIFuHD [29], ECON [41], SiTH [8]) and the baseline models introduced in Sec. 4.2 of the main paper. Specifically, Fig. 13 presents results of the methods not included in Fig. 6 of the main paper on synHOR dataset, while Fig. 14 illustrates reconstructed human-object shapes on the BEHAVE [2] dataset that were omitted from Fig. 7Fig. 15 and Fig. 16 show new examples of reconstructed shapes using the considered methods on synHOR and BEHAVE dataset, respectively. Across all the presented figures, our approach reconstructs the most realistic human-object shapes with the fewest artefacts compared to related works. While human-focused reconstruction approaches achieve higher detail in human shapes, they cannot reconstruct objects, failing to achieve our goal of joint human-object reconstruction. Even retraining PIFuHD on synHOR ($\mathrm{PIFuHD}_{\mathrm{ho}}$), incorporating pose features ($\sigma\mathrm{PIFuHD}_{\mathrm{ho}}$,) or using two MLPs ($2\mathrm{PIFuHD}_{\mathrm{ho}}$), does not sufficiently improve results, with severe depth ambiguities in the reconstructed objects. The proposed baselines that leverage pose features along with two MLPs ($2\sigma\mathrm{PIFuHD}_{\mathrm{ho}}$, $2\sigma\mathrm{PIFuHD}^{all}_{\mathrm{ho}}$, $2\sigma\mathrm{PIFuHD}^{sep}_{\mathrm{ho}}$) can reconstruct objects but are significantly more prone to noise and artifacts compared to SCHOR. The attention-based neural implicit model introduced by SCHOR ensures the joint reconstruction of spatially coherent and realistic human-object shapes from single images, outperforming all related approaches.
@@ 11 Discussion
Existing works [41, 8, 29] for high-quality human reconstruction are not designed to reconstruct objects. Consequently, when these methods are applied to real images containing both humans and objects, they behave differently in accordance with their network design, as illustrated in Fig. 14. For instance, if these methods are conditioned on the SMPL model [8, 41], the object is not reconstructed unless it is very close to the human. In such cases, these methods misinterpret the object as part of the clothing, merging it with the human mesh and resulting in a lack of 3D spatial coherence. Similarly, methods that integrate normals [29] to reconstruct human shapes struggle to distinguish between clothing and objects, treating them as a single entity.
One potential solution is to crop the object out of the image and reconstruct only the human with the above methods. However, this requires reconstructing the object separately, without considering its spatial relation with the human. In contrast, our approach is explicitly designed to jointly reason about humans and objects, distinguishing them as two distinct yet connected entities and reconstructing both in a spatially coherent manner.
We also highlight the importance of extracting features from each considered input and merging them using transformer encoders. This allows a better understanding of the global scene, ensuring 3D spatial coherence. In addition, learning local information about humans and objects allows our method to reproduce fine details in the human-object shapes.
Despite its strengths, our approach is prone to certain limitations. First, the reconstruction quality of smaller body parts, such as fingers and hair, can be improved. This refinement is essential for explicitly modelling human-object interactions and will be addressed in future work. Furthermore, the reconstruction quality of unseen regions in the full-body human image, such as the sides and back, is limited.
Another limitation arises from estimating the object pose features from known templates. This restricts the method’s ability to generalize to objects without predefined templates. Future work will explore strategies to estimate pose parameters without using templates.

### v2

@@ 9.3 Inpainting module
ReCHOR (ours) 14.23 17.15 .8902 .8241 48.84
Table 6: Quantitative evaluation of changing the diffusion module of ReCHOR. The human-object shapes reconstructed with the attention-based neural implicit module of ReCHOR are used for evaluation.
@@
In scenes containing both humans and objects, the object often occlude parts of the human body. When only partial regions of the body are visible, the network fails to reconstruct the occluded regions. This limitation is shown in Fig. 10, where ReCHOR is applied without the inpainting module to generate a full-body human image. Instead, the partial human image is directly input into the attention-based neural implicit model, which fails to reconstruct these occluded regions, producing holes in the human shapes. Inpainting the missing body regions in the input image is crucial to reconstruct realistic clothed human and object shapes.
ReCHOR addresses the challenge of human occlusion by leveraging the generative capability of diffusion models to inpaint the occluded body regions. A fine-tuning strategy is adopted to optimize the cross-attention layers of a pre-trained diffusion model, conditioning it on the mask of the body regions requiring inpainting. In this section, we explore different solutions for inpainting, including MAT [li2022mat], a transformer-based model for large hole inpainting, conditioning the diffusion model with the SMPL [25] normal map (Normal), and conditioning with both the mask and the normal map (Both). The effect of using these approaches for the inpainting module of ReCHOR is quantitatively evaluated by estimating the implicit representation of human-object shapes using their generated full-body human image (Tab. 6). Examples of human-object reconstructions using these methods for inpainting are shown in Fig. 12 while Fig. 11 illustrates the input images shown in the paper and generated using the discussed inpainting methods.
@@
Note that we do not quantitatively evaluate the generated images against ground truth, since the main goal of ReCHOR is 3D reconstruction rather than inpainting.
Figure 12: Human-object shapes reconstructed by changing the diffusion module of ReCHOR. Front view with and without the object and side views without the object are shown from synHOR in the top row, from BEHAVE [2] in the bottom.
@@ 10 Additional visual comparisons
In this section, we present additional visual results of human-object shapes reconstructed using ReCHOR, as well as comparisons with related methods that aim to reconstruct high-quality 3D humans (PIFuHD [27], ECON [39], SiTH [8]) and the baseline models introduced in Sec. 4.1 of the main paper. Specifically, Fig. 13 presents results of the methods not included in Fig. 5 of the main paper on synHOR dataset, while Fig. 14 illustrates reconstructed human-object shapes on the BEHAVE [2] dataset that were omitted from Fig. 4Fig. 15 and Fig. 16 show new examples of reconstructed shapes using the considered methods on synHOR and BEHAVE dataset, respectively. Across all the presented figures, our approach reconstructs the most realistic human-object shapes with the fewest artefacts compared to related works. Human-focused reconstruction approaches cannot reconstruct objects, failing to achieve our goal of joint human-object reconstruction. Even retraining PIFuHD on synHOR ($\mathrm{PIFuHD}_{\mathrm{ho}}$), incorporating pose features ($\sigma\mathrm{PIFuHD}_{\mathrm{ho}}$,) or using two MLPs ($2\mathrm{PIFuHD}_{\mathrm{ho}}$), does not sufficiently improve results, with severe depth ambiguities in the reconstructed objects. The proposed baselines that leverage pose features along with two MLPs ($2\sigma\mathrm{PIFuHD}_{\mathrm{ho}}$, $2\sigma\mathrm{PIFuHD}^{all}_{\mathrm{ho}}$, $2\sigma\mathrm{PIFuHD}^{sep}_{\mathrm{ho}}$) can reconstruct objects but are significantly more prone to noise and artifacts compared to ReCHOR. The attention-based neural implicit model introduced by ReCHOR ensures the joint reconstruction of realistic clothed human and object shapes from single images, outperforming all related approaches.
@@ 11 Discussion
Existing works [39, 8, 27] for realistic human reconstruction are not designed to reconstruct objects. Consequently, when these methods are applied to real images containing both humans and objects, they behave differently in accordance with their network design, as illustrated in Fig. 14. For instance, if these methods are conditioned on the SMPL model [8, 39], the object is not reconstructed unless it is very close to the human. In such cases, these methods misinterpret the object as part of the clothing, merging it with the human mesh and resulting in a severe depth ambiguity. Similarly, methods that integrate normals [27] to reconstruct human shapes struggle to distinguish between clothing and objects, treating them as a single entity.
One potential solution is to crop the object out of the image and reconstruct only the human with the above methods. However, this requires reconstructing the object separately, without considering its spatial relationship with the human. In contrast, our approach is explicitly designed to jointly reason about humans and objects, distinguishing them as two distinct yet connected entities and reconstructing both in a shared 3D space.
We also highlight the novelty of our feature extraction and fusion strategy, where pixel-aligned features from each input are merged via transformer encoders. This design allows the model to jointly learn both global and local contextual information about the scene. Capturing global information improves the understanding of human-object spatial relationships, while local information allows the reconstruction of realistic shapes.
Despite its strengths, our approach is prone to certain limitations. First, the reconstruction quality of smaller body parts, such as fingers and hair, can be improved. This refinement will improve the modeling of human-object interactions and will be addressed in future work. In addition, increasing the quality of the ground-truth object shapes of our dataset will allow the network to learn finer details for the reconstructed object shapes, further increasing the realism of the reconstruction.
Although ReCHOR can reconstruct arbitrary object shapes, it currently relies on object pose priors from a method [34] that requires known object templates, limiting its generalization to objects without predefined templates. This limitation can be addressed in the future by integrating priors from recent template-free methods [36].